OpenStudy (anonymous):

Does the the distance from (-1,7) to (-5,-4)= sqrt(137)? I'm not sure if I did it right.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Can it not be simplified any more?

OpenStudy (anonymous):

Nope. You can find it in decimals if you want to

OpenStudy (anonymous):

I need it in the form of a root. \[\sqrt{137}\] is the final answer?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

\[\sqrt{(-1+5)^2+(7+4)^2}=\sqrt{4^2+11^2}=\sqrt{16+121}=\sqrt{137}\] just practicing my latex

OpenStudy (anonymous):

Using the same formula, then the distance from (-4,-6) to (4,-5) should be \[\sqrt{65}\] correct?

OpenStudy (anonymous):

yes. distance from -4 to 4 is 8. distance from -6 to -5 is 1 \[8^2+1^2=65\]

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

one last question. \[\sqrt{8}\]=\[2\sqrt{2}\] and not just 2 right?